The population P of a city if given by
P=2500e^(kt)
where t is time of the years , with t=0 corresponding to the years 1980 .In 1935 ,the population was 3350 .Find the value of k and use this result to predict the population in the years 2000

Answer by tommyt3rd(5050)

t=1935-1980= -45
3350=2500e^(k*(-45))

e^(k*(-45))=3350/2500
ln(e^(k*(-45)))=ln(3350/2500)

k=ln(3350/2500)/(-45)=-0.0065
In 2000 t=20, so...
P=2500e^(-0.0065*20)= 2196

:)